Your grandmother has set up an annuity of $4000 in an account that pays 5.2%/a compounded monthly. What equal monthly payments will the annuity provide for in the next 4 years?


1
Expert's answer
2021-02-24T06:42:56-0500

"Fut""ure value =C {(1+r\/k)"nk -1}/(r/k)

Fv = Future value = $4000

r = interest rate = 5.2%

n = years = 4

k = 12

C = monthly instalment = ?

4000 = C{(1+0.025/12)4X12 - 1}/(0.052/12)

4000 = C{(1.004333)48 - 1}/(0.004333)

4000 = C(1.23064 - 1)/(0.004333)

4000 = C(0.23064/0.004333)

4000 = C(53.2288)

C = 4000/53.2288

C = 75.15

Monthly payments = $75.15